OpenStudy (anonymous):

Could somone help me with this one Algebra1 problem? 8 + x - 1 over x + 4 Thank you!!!!

OpenStudy (anonymous):

\[\frac{8+x-1}{x+4}=\frac{7+x}{x+4}\] not much else you can do with it

OpenStudy (anonymous):

the eight is by itself its next to the x-1 and thank you

OpenStudy (anonymous):

\[8+\frac{x-1}{x+4}=\frac{8(x+4)+x-1}{x+4}=\frac{8x+32+x-1}{x+4}=\frac{9x+31}{x+4}\]

OpenStudy (anonymous):

yeah that the right one

OpenStudy (anonymous):

hope the steps are clear

OpenStudy (anonymous):

They are Thank you @satellite73